# The State of Open Science in Software Engineering Research: A Case Study of ICSE Artifacts

### v2

@@ 3.2.3. Installing Dependencies and External Tools
$\bullet$ Inferring Dependencies Without Documentation. In many cases, artifacts lacked both installation instructions and explicit dependency lists. When this occurred, we searched for language-specific dependency files within the artifact. For Python artifacts, we installed Python and its package manager, then looked for files like requirements.txt or environment.yml. We installed OpenJDK and Maven for Java artifacts and searched for files such as pom.xml. If no such files were present, we manually inspected import statements and installed dependencies one by one, a time-consuming and error-prone process. Unresolved dependencies at this stage were deferred to the troubleshooting phase. When version information was missing, we applied the two-year backward buffer heuristic, similar to the previous step, to estimate suitable versions.
@@ 3.2.4. Running the Experiment
After installing all dependencies, we proceeded to execute the artifact. We first consulted the documentation to identify any executable commands provided by the authors. The type and granularity of instructions varied across artifacts. For example, studies involving Large Language Models (LLMs) often included commands for pretraining, fine-tuning, and evaluation. Others provided shell scripts to trigger an end-to-end execution pipeline or step-by-step instructions to run individual components.
However, many artifacts lacked execution instructions entirely. In such cases, we inspected the codebase to infer a plausible execution entry point. The success of this process depended heavily on the technology stack and project structure. For instance, executing a standalone Python script is typically straightforward, unless undocumented runtime parameters are required. In more complex cases involving multiple scripts or loosely organized codebases, identifying the correct execution sequence without guidance was challenging. Some programming languages follow standard convention regarding entry-points (e.g., main.java in Java projects), while others rely on user-defined conventions such as naming scripts main.py, run.py, or run.sh. These cues often helped us infer where to begin, though no consistent standard was observed. In such cases, we relied on intuition, programming experience, and a structural understanding of the codebase.
Once a command was executed, we monitored the outcome. If it completed successfully, we proceeded to the next step. If the process failed with a runtime error, we deferred to the troubleshooting phase.
@@
M3 Ubuntu 22.04.5 LTS, 64 bit Intel® Core™ i7-11700U @ 2.6 GHz 32 3170 GB 1 TB 3 * NVIDIA A100 80 GB PCIe
@@
M5 Windows 11, 64 bit Intel(R) Xeon(R) Platinum 8356H @ 3.90 GHz 4 32 GB 1 TB Intel® Iris® Xe
@@ 3.2.5. Troubleshooting and Fixing Runtime Errors
(ii) Modify Instructions. In many cases, we modified the provided instructions, for example, adjusting relative paths, correcting flag values, or adding missing commands that were not included in the documentation. For example, we changed the -nproc_per_node flag value to adapt to the number of GPUs in our environment.
(iii) Source Code Modification. To resolve runtime errors, sometimes we modify the artifact’s source code. This was often challenging, as poor or missing code documentation made it difficult to infer the intended logic and behavior (e.g., changing the input directory path in the source code to access the data file).
(iv) Change File Organization. In many cases, the execution failed due to the wrong file structure, missing files, or directories. We addressed these by reorganizing folders, creating missing directories, removing existing directories, or reconstructing essential files needed for execution.
(v) Adjust Configuration Files. We updated configuration files to align with our environment, including fixing directory paths and modifying control variables. For example, we changed the dataset path in config.json to enable the successful execution of an artifact.
@@ 3.2.6. Labeling
$\bullet$ Executable – able to complete the execution with or without modifications.
$\bullet$ Partially Executable – able to execute some component of the artifact with or without modification.
$\bullet$ Not Executable – can not execute any component of the artifact following the given instructions and troubleshooting.
We further classified each of the above levels based on effort level and modification required. Effort level is defined based on the total time spent attempting to execute the artifact: Low Effort – less than 2 hours, Moderate Effort – between 2 and 8 hours, and High Effort – more than 8 hours. Modification level reflects the time invested in troubleshooting and fixing issues before determining the final execution status: Minor Modification – less than 1 hour, Major Modification – between 1 and 4 hours, and Critical Modification – more than 4 hours. The challenges and time required to complete the execution process were documented for further qualitative analysis.
@@ 3.4. Examining Reproducibility of Executable Artifacts and the Factors Behind Failure
We assessed reproducibility of results through a three-step process for all artifacts labeled as Executable or Partially Executable. First, we documented the outputs generated during execution, whether partial or complete. Second, we checked whether the artifact included explicit validation instructions to compare the generated outputs with those reported in the original study. These instructions typically included one or more of the following: (i) reference outputs or logs, (ii) scripts or commands to compute performance metrics, or (iii) visualizations or result plots for comparison. When such validation steps were available, we followed them to verify whether our outputs matched or closely resembled the original results. Finally, without validation guidance, we manually compared our outputs against results reported in the paper, focusing on the described metrics, tables, and figures. However, when the execution produced raw outputs that were not directly represented in the paper, and no scripts were provided to process or interpret them, we could not determine reproducibility. Based on the outcome of this evaluation, we assigned each artifact to one of five reproducibility categories:
$\bullet$ Fully Reproducible – the artifact executed successfully and the results matched those reported in the original paper using either provided validation steps or manual comparison.
$\bullet$ Partially Reproducible – the artifact executed successfully or partially, but only a subset of the outputs matched those reported in the original study.
$\bullet$ Not Reproducible – the artifact executed successfully or partially, but none of the outputs matched the original study results, even after manual comparison.
$\bullet$ Unverifiable – the artifact executed successfully or partially, but verification was not feasible due to missing validation instructions, reference results, or guidance on how to interpret the results.
$\bullet$ No Output Generated – the artifact executed successfully or partially but produced no output.
This systematic evaluation provides one of the first large-scale empirical assessments of whether ICSE artifacts can actually reproduce the findings reported in their corresponding papers.
@@ 3.5. Experimental Setup
To conduct the execution and reproducibility experiments, the first two authors analyzed artifacts using a total of seven machines, selected to closely align with the requirements specified in the replication packages. Table 2 provides an overview of the hardware and operating system specifications used in this study. Each evaluator had access to at least one Windows machine for running artifacts that provided Windows-specific instructions or executables, and two Ubuntu-based Linux machines, at least one of which was equipped with a GPU to support experiments involving hardware acceleration or deep learning workloads.
We kept a Mac machine as a backup and needed it only for one artifact that required Mac-specific data processing.
@@ 4. Study Findings
This section presents our empirical findings in response to the four research questions, supported by data and evidence from our analysis.
@@ 4.1. Answering RQ1: Executability Status, Level of Effort and Modifications to Execute Replication Package
Executability of Replication Package. Figure 3 shows the executability status of artifacts over the years. In 2024, only 7 out of 22 artifacts (31.82%) were fully executable, with 8 (36.36%) not executable and 7 (31.82%) partially executable. A similar pattern appears in 2023, where 10 out of 20 artifacts (50%) executed successfully, but 6 (30%) failed completely, and 4 (20%) were only partially runnable. These figures suggest that recent artifacts are not consistently more executable despite increased availability. The most promising year was 2021, with 8 out of 11 artifacts (72.73%) fully executable, indicating a potential high point in artifact quality. However, this trend was not sustained in later years. Earlier years (2015–2018) had very few artifacts overall, limiting inference, but notably, all artifacts in 2015 and 2016 failed to execute.
@@
In total, across all years, 40 out of 100 artifacts (40%) were executable, 24 (24%) partially executable, and 36 (36%) not executable. These results reveal that while artifact sharing has increased, a significant fraction still requires substantial improvement to meet execution standards. The data underscore that growth in replication artifacts does not automatically translate to better executability. Without targeted community efforts on environment specification, documentation, and packaging standards, the reproducibility goals of open science will remain elusive.
Level of Effort Required to Execute Replication Packages. Figure 4(a) presents the relationship between executability status and the level of effort required to execute the artifacts. The results show a stark contrast in effort distribution across executable, partially executable, and non-executable artifacts. Among the 40 executable artifacts, the majority (72.5%) required either low (11) or moderate (18) effort. However, 11 executable artifacts (27.5%) still demanded high effort, indicating that even runnable artifacts often lack sufficient documentation, packaging, or automation support. The situation is more difficult for partially executable artifacts (24). All required moderate (13) or high (11) effort, yet none could be executed successfully. Even for the non-executable artifacts (36), we invested non-trivial time: 52.78% (19/36) required moderate to high effort before reaching a dead end. The remaining 47.22% (17/36) failed early due to severe deficiencies or incompleteness and were classified under low effort, not because they were simple, but because they were unworkable from the start. These results emphasize the urgent need for community-wide practices that minimize unnecessary overhead by promoting thorough environment specification, proper automation, and quality assurance of replication packages.
Figure 5. Types of Modifications
@@
Level of Modification Required to Execute Replication Packages. Figure 4(b) visualizes the extent of modifications needed across artifacts with varying executability statuses. Among the 40 executable artifacts, only 13 (32.5%) worked without any modification. An additional 6 artifacts (15%) required only minor changes, such as adjusting directory paths or fixing broken shell commands. However, a significant portion of the executable artifacts (37.5%) demanded major to critical modifications, including editing source code or resolving complex environment conflicts, just to achieve successful execution. Such findings challenge the assumption that “executable” implies low effort.
The situation deteriorates for partially executable artifacts. Of the 24 analyzed, 79.17% (19/24) required major or critical modifications. These typically involved deep architectural issues, undocumented assumptions, or persistent runtime errors that made full reproduction infeasible despite extensive manual intervention. For non-executable artifacts (36), 20 (55.56%) reached major or critical modification levels before ultimately failing. This indicates that significant effort was still invested in troubleshooting, debugging, and altering code, but the artifacts could not be salvaged. The remaining 16 artifacts (44.44%) required no or minor modification, not because they were simple, but because execution attempts were abandoned early due to insurmountable issues such as missing datasets or unusable documentation.
These results reaffirm that sharing alone is not enough. Without proper environmental details, dependency control, and clear code, artifacts often remain unusable. The high modification burden highlights a pressing need for stronger tooling, standards, and author-side validation.
@@ 4.2. Answering RQ2: Types of Modifications Needed to Execute Replication Packages
Figure 5 shows that environment-related issues were most prevalent, affecting 57% of artifacts. Common problems included installing missing dependencies (43.86%) and resolving version conflicts (45.61%), often due to unpinned dependencies or outdated instructions. In one case, resolving a circular Python dependency took nearly 10 hours–avoidable with proper versioning. Instruction-related changes appeared in 31 artifacts, where 92.86% of execution commands required adjustments such as fixing syntax, flags, or paths. A notable case involved a five-hour debugging effort caused by a missing –gpus flag in a Docker command, leading to a misleading “division by zero” error.
Code-level modifications were needed in 22% of the artifacts, mainly to fix runtime errors (90.91%). These were especially challenging due to poor code documentation, as only 26.67% (4/15) of such artifacts were ultimately executed, often with high effort. One required modifying the tensor2tensor library logic twice, highlighting a broader lack of maintainability. The remaining 22% of artifacts required surface-level changes. Specifically, 16% needed folder restructuring to fix broken paths, and 6% required environment variable or path adjustments in config files. Though simple, such issues were common and pointed to inadequate quality control in artifact preparation.
Figure 6. List of Challenges
A hierarchical chart displaying execution barriers organized into 3 main themes: Documentation Issues, Environment Setup Challenges, and Technical Issues. Each theme branches into specific challenge types with frequency counts.
@@ 4.3. Answering RQ3: Challenges that Prevent Executability of Replication Packages
Similar to the modification patterns, environment setup challenges emerged as the most frequent issue, affecting 91.67% of the artifacts with execution failures. Within this category, 85.45% of the artifacts did not provide essential system specifications. These omissions included missing operating system details, platform-specific configurations, or hardware-related requirements such as GPU support, RAM, or disk space. In 60% of the artifacts, dependency information such as required libraries or external tools were not documented. Furthermore, 54.55% failed to mention the versions of the programming languages or dependencies used in the project (e.g., Python, Java), making it difficult to align our environment with that of the original experiment. While any one of these issues alone might not entirely prevent execution, their combination created significant barriers. For instance, 35% of the artifacts were still executable when only one of these issues was present. However, for artifacts that suffered from all three challenges simultaneously, the executability rate dropped sharply to just 8%, with more than half (52%) of them failing to run altogether. In contrast, when none of these three environment-related issues were present, 78.8% of the artifacts executed successfully, underscoring the importance of providing clear and complete environment setup instructions in replication packages.
@@
The second major challenge category pertained to inadequate documentation, which affected 70% of the non or partially executable artifacts. Within this group, we observed three recurring problems. First, in 57.14% of the artifacts, the documentation was ambiguous where steps were either unclear or not logically ordered, making it difficult to follow. Second, 45.24% of the artifacts had missing instructions, with key steps in the execution process entirely omitted. While we were sometimes able to overcome these issues using our technical expertise and by inferring missing details from the provided materials, success was inconsistent. When only one of these documentation issues was present, we could still execute 42.4% of the artifacts. However, when both ambiguous and missing documentation occurred together, the executability rate dropped to 7.7%, with 61.5% of those cases ending in partial execution. The situation was especially severe when artifacts lacked any usable documentation at all. In this subset categorized as ”No Documentation”, we were able to execute only 7.1% of the artifacts, while the remaining 92.9% could not be run at all.
@@
The final set of challenges related to inaccessible or outdated components in the replication packages. For artifacts with missing project elements such as datasets, scripts, or configuration files, only 15% could be executed. Success in these few cases was possible either because the authors had provided a Docker image alongside the code base or because we were able to reconstruct missing files using clues from other resources. Outdated components were another common obstacle. In these cases, changes had been made to parts of the repository while other dependent components were left untouched, causing inconsistency and execution errors. Only 18.2% of these artifacts were executable, while the majority (63.6%) could only be partially executed due to unresolved conflicts or deprecated configurations.
@@
These findings reveal that while many artifacts are shared with the intention of supporting reuse, a large proportion fall short due to avoidable issues such as incomplete documentation, missing configuration details, or outdated codebases. Addressing these challenges through clearer documentation, robust environment specifications, and well-maintained artifact repositories would significantly improve the executability of shared artifacts and move the SE community closer to the goals of open science.

### v4

@@ 3.2.3. Installing Dependencies and External Tools
$\bullet$ Inferring Dependencies Without Documentation. In many cases, artifacts lacked both installation instructions and explicit dependency lists. In these cases, we searched the artifact for language-specific dependency files. For Python artifacts, we installed Python and its package manager, then searched for files such as require-
ments.txt
or environment.yml. We installed OpenJDK and Maven for Java artifacts and searched for files such as pom.xml. If no such files were present, we manually inspected import statements and installed dependencies one by one, a time-consuming and error-prone process. Unresolved dependencies at this stage were deferred to the troubleshooting phase. When version information was missing, we applied the two-year backward buffer heuristic, similar to the previous step, to estimate suitable versions.
@@ 3.2.4. Running the Experiment
After installing all dependencies, we attempted to execute the artifact. We first consulted the documentation to identify any executable commands provided by the authors. The type and granularity of instructions varied across artifacts. For example, studies involving Large Language Models often included commands for pretraining, fine-tuning, and evaluation. Others provided shell scripts to trigger an end-to-end execution pipeline or step-by-step instructions to run individual components.
However, many artifacts lacked execution instructions entirely. In such cases, we inspected the codebase to infer a plausible entry point for execution. The success of this process depended heavily on the technology stack and project structure. For instance, executing a standalone Python script is typically straightforward, unless undocumented runtime parameters are required. In more complex cases involving multiple scripts or loosely organized codebases, identifying the correct execution sequence without guidance was challenging. Some programming languages follow standard conventions for entry points (e.g., main.java in Java projects), whereas others rely on user-defined conventions, such as naming scripts main.py, run.py, or run.sh. These cues often helped us infer where to begin, though no consistent standard was observed. In such cases, we relied on intuition, programming experience, and a structural understanding of the codebase.
Once a command was executed, we monitored the outcome. If it completed successfully, we proceeded to the next step. If it failed with a runtime error, we deferred to the troubleshooting phase.
@@
M3 Ubuntu 22.04.5 LTS, 64 bit Intel(R) Xeon(R) Platinum 8356H CPU @ 3.90 GHz 32 3 TB 24 TB 3 * NVIDIA A100 80 GB PCIe
@@
M5 Windows 11, 64 bit Intel® Core™ i7-11700U @ 2.6 GHz 4 32 GB 1 TB Intel® Iris® Xe
@@ 3.2.5. Troubleshooting and Fixing Runtime Errors
(ii) Modify Instructions. In many cases, we modified the provided instructions by adjusting relative paths, correcting flag values, or adding missing commands that were not included in the documentation. For example, we changed the value of -nproc_per_node flag to match the number of GPUs in our environment.
(iii) Source Code Modification. To resolve runtime errors, sometimes we made modifications to the artifact’s source code. This was often challenging, as poor or missing code documentation made it difficult to infer the intended logic and behavior (e.g., changing the input directory path in the source code to access the data file).
(iv) Change File Organization. In many cases, execution failed due to incorrect file structure, missing files, or directories. We addressed these issues by reorganizing folders, creating missing directories, removing unnecessary directories, or reconstructing essential files required for execution.
(v) Adjust Configuration Files. We updated configuration files to align with our environment, including fixing directory paths and modifying control variables. For example, we updated the dataset path in config.json to enable the successful execution of the artifact.
@@ 3.2.6. Labeling
  • Executable – able to complete the execution with or without modifications.
  • Partially Executable – able to execute some component of the artifact with or without modification.
  • Not Executable – cannot execute any component of the artifact following the given instructions and troubleshooting.
We further classified each of the above levels by effort level and required modification. Effort level is defined based on the total time spent attempting to execute the artifact: Low Effort – less than 2 hours, Moderate Effort – between 2 and 8 hours, and High Effort – more than 8 hours. Modification level reflects the time invested in troubleshooting and fixing issues before determining the final execution status: Minor Modification – less than 1 hour, Major Modification – between 1 and 4 hours, and Critical Modification – more than 4 hours. The challenges we encountered and the time required to complete the execution process were documented for further qualitative analysis.
@@ 3.4. Examining Reproducibility of Executable Artifacts and the Factors Behind Failure
We assessed reproducibility of results through a three-step process for all artifacts labeled as Executable or Partially Executable. First, we documented the outputs generated during execution, whether partial or complete. Second, we checked whether the artifact included explicit validation instructions to compare the generated outputs with those reported in the original study. These instructions typically included one or more of the following: (i) reference outputs or logs, (ii) scripts or commands to compute performance metrics, or (iii) visualizations or result plots for comparison. When such validation steps were available, we followed them to verify whether our outputs matched or closely resembled the original results. Finally, without validation guidance, we manually compared our outputs with results reported in the paper, focusing on the metrics, tables, and figures described. However, when the execution produced raw outputs that were not directly reported in the paper, and no scripts were provided to process or interpret them, we could not assess reproducibility. Based on the outcome of this evaluation, we assigned each artifact to one of five reproducibility categories:
  • Fully Reproducible – the artifact executed successfully and the results matched those reported in the original paper using either provided validation steps or manual comparison.
  • Partially Reproducible – the artifact executed successfully or partially, but only a subset of the outputs matched those reported in the original study.
  • Not Reproducible – the artifact executed successfully or partially, but none of the outputs matched the original study results, even after manual comparison.
  • Unverifiable – the artifact executed successfully or partially, but verification was not feasible due to missing validation instructions, reference results, or guidance on how to interpret the results.
  • No Output Generated – the artifact executed successfully or partially but produced no output.
This systematic evaluation provides a large-scale empirical assessments of whether ICSE artifacts can actually reproduce the findings reported in their corresponding papers.
@@ 3.5. Experimental Setup
To conduct the execution and reproducibility experiments, the first two authors analyzed artifacts using a total of seven machines, selected to closely align with the requirements specified in the replication packages. Table 2 provides an overview of the hardware and operating system specifications used in this study. Each evaluator had access to at least one Windows machine to run artifacts that provided Windows-specific instructions or executables, and to two Ubuntu-based Linux machines, at least one of which was equipped with a GPU to support experiments involving hardware acceleration or deep learning workloads. We kept a Mac machine as a backup, and it was needed only for one artifact that required Mac-specific data processing.
@@ 4. Study Findings
This section presents the empirical findings for our four research questions, supported by data and evidence from our analysis.
@@ 4.1. Answering RQ1: Executability Status, Level of Effort and Modifications to Execute Replication Package
Executability of Replication Package. Figure 3 shows the executability status of artifacts over the years. In 2024, only 7 of 22 artifacts (31.82%) were fully executable, 8 (36.36%) were not executable, and 7 (31.82%) were partially executable. In contrast, 2023 shows a higher rate of full executability, with 10 of 20 artifacts (50%) executing successfully, while 6 (30%) failed completely and 4 (20%) achieved only partial execution. These figures suggest that recent artifacts are not consistently more executable despite increased availability. The most promising year was 2021, with 8 out of 11 artifacts (72.73%) fully executable, indicating a potential high point in artifact quality. However, this trend was not sustained in later years. In earlier years (2015-2018), there were very few artifacts overall, which limited inference. Notably, all artifacts in 2015 and 2016 failed to execute.
@@
In total, across all years, 40 out of 100 artifacts (40%) were executable, 24 (24%) partially executable, and 36 (36%) not executable. These results reveal that while artifact sharing has increased, a significant fraction still requires substantial improvement to meet execution standards. The data suggests that growth in artifact sharing does not necessarily translate to improved executability. Without targeted community efforts to specify the environment, document, and establish packaging standards, the executability goals of open science will remain elusive.
Level of Effort Required to Execute Replication Packages. Figure 4(a) presents the relationship between executability status and the level of effort required to execute the artifacts. The results show a stark contrast in effort distribution across executable, partially executable, and non-executable artifacts. Among the 40 executable artifacts, the majority (72.5%) required either low (11) or moderate (18) effort. However, 11 executable artifacts (27.5%) still demanded high effort, indicating that even runnable artifacts often lack sufficient documentation, packaging, or automation support. The situation is more difficult for partially executable artifacts. Among the 24 partially executable artifacts, 13 required moderate, and 11 required high effort, yet none could be executed successfully. Even for the non-executable artifacts (36), we invested non-trivial time: 52.78% (19/36) required moderate-to-high effort before reaching a dead end. The remaining 47.22% (17/36) failed early due to severe deficiencies or incompleteness and were classified under low effort, not because they were simple, but because they were unworkable from the start. These results emphasize the urgent need for community-wide practices that minimize unnecessary overhead by promoting thorough environment specification, proper automation, and quality assurance of replication packages.
Figure 5. Types of Modifications Required for Execution
@@
Level of Modification Required to Execute Replication Packages. Figure 4(b) visualizes the extent of modifications needed across artifacts with varying executability statuses. Among the 40 executable artifacts, only 13 (32.5%) worked without any modification. An additional 6 artifacts (15%) required only minor changes, such as adjusting directory paths or fixing broken shell commands. However, a significant portion of the executable artifacts (37.5%) required major-to-critical modifications, including editing source code or resolving complex environment conflicts, to achieve successful execution. Such findings challenge the assumption that “executable” implies low effort.
The situation deteriorates for partially executable artifacts. Of the 24 analyzed, 79.17% (19/24) required major or critical modifications. These typically involved deep architectural issues, undocumented assumptions, or persistent runtime errors, making full reproduction infeasible despite extensive manual intervention. Among 36 non-executable artifacts, 20 (55.56%) underwent major or critical modifications before ultimately failing. This indicates that significant effort was still invested in troubleshooting, debugging, and altering code, but the artifacts could not be salvaged. The remaining 16 artifacts (44.44%) required no or minor modification, not because they were simple, but because execution attempts were abandoned early due to insurmountable issues such as missing datasets or unusable documentation.
Overall, these results suggest that artifact sharing alone is insufficient. Without proper environmental details, dependency control, and clear code, artifacts often remain unusable. The high modification burden highlights a pressing need for stronger tooling, standards, and author-side validation.
@@ 4.2. Answering RQ2: Types of Modifications Needed to Execute Replication Packages
Figure 5 shows that environment-related issues were most prevalent, affecting 57% of artifacts. Common problems included installing missing dependencies (43.86%) and resolving version conflicts (45.61%), often due to unpinned dependencies or outdated instructions. In one case, resolving a circular Python dependency took nearly 10 hours, which could be avoided with proper versioning. Instruction-related changes occurred in 31 artifacts, with 92.86% of execution commands requiring adjustments, such as fixing syntax, flags, or paths. A notable case involved a five-hour debugging effort due to a missing –gpus flag in a Docker command, resulting in a misleading “division by zero” error.
Code-level modifications were required in 22% of the artifacts, primarily to address runtime errors (90.91%). These were particularly challenging due to poor code documentation, as only 26.67% (4/15) of such artifacts were ultimately executed, often requiring substantial effort. One required modifying the tensor2tensor library logic twice, highlighting a broader lack of maintainability. The remaining 22% of artifacts required surface-level changes. Specifically, 16% needed folder restructuring to fix broken paths, and 6% required environment variable or path adjustments in config files. Although simple, such issues were common and indicated inadequate quality control in artifact preparation.
Figure 6. List of Challenges Preventing Execution
A hierarchical chart displaying execution barriers organized into 3 main themes: Documentation Issues, Environment Setup Challenges, and Technical Issues. Each theme branches into specific challenge types, with frequency counts provided.
@@ 4.3. Answering RQ3: Challenges that Prevent Executability of Replication Packages
Similar to the modification patterns, environment setup challenges emerged as the most frequent issue, affecting 91.67% of the artifacts with execution failures. Within this category, 85.45% of the artifacts lacked essential system specifications. These omissions included missing operating system details, platform-specific configurations, or hardware-related requirements such as GPU support, RAM, or disk space. In 60% of the artifacts, dependency information, such as required libraries or external tools, was not documented. Furthermore, 54.55% failed to mention the versions of the programming languages or dependencies used in the project (e.g., Python, Java), making it difficult to align our environment with that of the original experiment. While any one of these issues alone might not entirely prevent execution, their combination created significant barriers. For instance, 35% of the artifacts were still executable when only one of these issues was present. However, for artifacts that faced all three challenges simultaneously, the executability rate dropped sharply to 8%, with more than half (52%) failing to run. In contrast, when none of these three environment-related issues were present, 78.8% of the artifacts executed successfully, underscoring the importance of providing clear and complete environment setup instructions in replication packages.
@@
The second major challenge category was inadequate documentation, affecting 70% of non-executable or partially executable artifacts. Within this group, we observed three recurring problems. First, in 57.1% of the artifacts, the documentation was ambiguous, where steps were either unclear or not logically ordered, making it difficult to follow. Second, 45.2% of the artifacts had missing instructions, with key steps in the execution process entirely omitted. While we were sometimes able to overcome these issues using our technical expertise and by inferring missing details from the provided materials, success was inconsistent. When only one of these documentation issues was present, we could still execute 42.4% of the artifacts. However, when both ambiguous and missing documentation occurred together, the executability rate dropped to 7.7%, with 61.5% of those cases ending in partial execution. The situation was especially difficult when artifacts lacked any usable documentation. In this subset categorized as ”No Documentation”, we were able to execute only 7.1% of the artifacts, whereas the remaining 92.9% could not be executed.
@@
The final set of challenges related to inaccessible or outdated components in the replication packages. For artifacts with missing project elements such as datasets, scripts, or configuration files, only 15% could be executed. Success in these few cases was achieved either because the authors provided a Docker image alongside the codebase or because we reconstructed missing files using clues from other resources. Outdated components were another common obstacle. In these cases, changes were made to parts of the repository while other dependent components were left untouched, leading to inconsistencies and execution errors. Only 18.2% of these artifacts were executable, while the majority (63.6%) could only be partially executed due to unresolved conflicts or deprecated configurations.
@@
These findings indicate that while many artifacts are shared with the intention of supporting reuse, a large proportion fall short due to avoidable issues such as incomplete documentation, missing configuration details, or outdated codebases. Addressing these challenges through clearer documentation, robust environment specifications, and well-maintained artifact repositories would significantly improve the executability of shared artifacts and move the SE community closer to the goals of open science.